z(x + y) p+z(x -y)q=x^2+y^2 , where p=dz/dx ,q=dz/dy
1
Expert's answer
2020-12-20T18:21:50-0500

P=z(x+y)=zx+yz

Q=z(x-y)=zx-zy

R=x2+y2

"\frac{dx}{zx+yz}=\frac{dy}{zx-zy}=\frac{dz}{x^2+y^2}"

Multiplyers x,-y,-z

xdx-ydy-zdz=0

1/2(x2-y2-z2)=C1

Multiplyers -x,y,z

-xdx+ydy+zdz=0

1/2(-x2+y2+z2)=C2